OpenStudy (anonymous):

Antonio is riding on a train. He gets up from his seat and walks forward. As he walks, he travels 1.1 miles in 2 minutes. As he walks back to his seat, he travels 0.9 miles in 2 minutes. How fast was Antonio walking inside the train?

OpenStudy (anonymous):

average speed = total distance tarveled / time taken = 1.1+0.9 / 2+2 = 2/4 = 1/2 = 0.5 miles/min

OpenStudy (anonymous):

*traveled